Three cells of emf 1.5v each and internal resistance 0.5ohms are connected in series to a 1.5ohms external resistor. Calculate the current flowing through the circuit.
1
Expert's answer
2020-02-28T10:19:31-0500

"I=\frac{E} {(R+R_{internal} )} =\frac{1.5 \times 3} {(1. 5+0.5×3)} =1.5 A"